You find the following Treasury bond quotes. To calculate the number of years until maturity, assume that it is currently May 2016. The bonds have a par value of $1,000.

Rate Maturity
Mo/Yr
Bid Asked Chg Ask
Yld
?? May 25 103.5501 103.5379 +.3054 6.059
5.374 May 30 104.4991 104.6448 +.4329 ??
6.178 May 40 ?? ?? +.5444 4.091

In the above table, find the Treasury bond that matures in May 2040. What is the asked price of this bond in dollars? (Do not round intermediate calculations and round your answer to 2 decimal places, e.g., 32.16.)

Asked price            $

If the bid-ask spread for this bond is .0545, what is the bid price in dollars? (Do not round intermediate calculations and round your answer to 2 decimal places, e.g., 32.16.)

Bid price            $
Answer #1

N = 24 * 2 = 48

I = 4.091%/ 2 = 2.0455%

PMT = 6.178% * 1,000/ 2

PMT = 30.89

FV = 1,000

Using Financial Calculator:

PV = 1,317.13

Part B:

Asked price = 1,317.13

Bid Price = 1,317.13 - (0.0545 * 1,000/100)

Bid Price = 1,316.59